OpenStudy (anonymous):

What is the sum of 12 – 5i and –3 + 4i? A) –16 + 63i B) 9 – i C) 9 – 9i D) 15 – 9i

OpenStudy (anonymous):

just add everything

OpenStudy (anonymous):

B

OpenStudy (anonymous):

how do u add i

OpenStudy (mathmale):

@Tannercollins : I'd suggest you align the real and imaginary terms in columns and then add up each column: 12 – 5i –3 + 4i ------- 12-3 = 9 (the real part of this expression) -5i+4i=(-5+4)i = -i (the imaginary part) Note that " i " represents the imaginary operator, and that \[i=\sqrt{-1}\]